# Just Ramp-up: Unleash the Potential of Regression-based Estimator for A/B Tests under Network Interference

### v2

@@ 5.2 Basic Simulation under Linear Interference
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+\beta_{2}\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\operatorname{deg}_{i}}+\epsilon_{i}$ (41)
where $\epsilon_{i}\sim N(0,\sigma_{e}^{2})$ are i.i.d. noises. We set the parameters as $(\beta_{0},\beta_{1},\beta_{2},\sigma_{e})=(1,1,1,0.1)$, and the true GATE is $\beta_{1}+\beta_{2}=2$.
@@
Furthermore, if cluster-level randomization is employed and the ramp-up process has reached the stage of $c_{t}=0.05$, the bias is only $-0.032$, which is two orders of magnitude smaller than the signal scale, i.e., $\beta_{1}$ and $\beta_{2}$. Although the standard deviation remains $0.121$, which is still notable, the GATE estimate at this stage can be considered credible. Finally, for $c_{t}=0.5$, the performance reaches a state-of-the-art level compared to settings where only base levels and outcomes under $c_{T}$ are available, as far as we know.
@@ 5.3 Multiplicative Potential Outcome Model
$Y_{i}(\mathbf{z})=(\beta_{0}+\epsilon_{i})\cdot\frac{\deg_{i}}{\overline{\deg}}\cdot(1+\beta_{1}z_{i}+\beta_{2}\frac{\sum_{j\in\mathcal{N}_{i}}z_{j}}{\deg_{i}})$ (44)
Here, $\overline{\deg}$ is the average degree among units, and $\epsilon\sim N(0,\sigma_{e}^{2})$ are i.i.d. noises. Similar to previous section, we set $(\beta_{0},\beta_{1},\beta_{2},\sigma_{e})=(1,1,1,0.1)$, and the true GATE is $\beta_{0}(\beta_{1}+\beta_{2})=2$. We first present the result in Table 9.
@@ 5.4 Quadratic and Square Root Interference Term
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+\beta_{2}\left(\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\operatorname{deg}_{i}}\right)^{2}+\epsilon_{i}$ (45)
The quadratic interference term indicates that the marginal benefit of allocating more treatments increases constantly, making extrapolation more challenging. We set $(\beta_{0},\beta_{1},\beta_{2},\sigma_{e})=(1,1,1,0.1)$, so the true GATE remains $\beta_{1}+\beta_{2}=2$. The results for our default setting are presented in Table 11, and the results with 2-step merged data, where $c_{1}=0$ and $c_{T}$ is varied, are shown in Table 12.
@@
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+\beta_{2}\sqrt{\left(\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\operatorname{deg}_{i}}\right)}+\epsilon_{i}$ (46)
with the same parameters $(\beta_{0},\beta_{1},\beta_{2},\sigma_{e})=(1,1,1,0.1)$. We present corresponding results in Table 13 and 14.
@@ 5.5 Multihop interference
Moreover, we implement the multihop interference as shown in Example 2. Specifically, we maintain the setting that $(\beta_{0},\beta_{1},\beta_{2},\sigma_{e})=(1,1,1,0.1)$, and set $M=2$ and $(r_{1},r_{2})=(1,0.5)$. Therefore, the true GATE in this case is given by:
$\text{GATE}=(\beta_{0}+\beta_{1}+\frac{\beta_{2}}{n}\mathbf{1}^{\top}B\mathbf{1})-\beta_{0}\approx 2.489$ (47)
The numerical outcome depends on the concrete network topology since we remove the circular effect that is introduced by the power of the (normalized) adjacency matrix. If we did not remove the diagonal terms of the interference matrix $B$, the GATE would be $\beta_{1}+\beta_{2}(r_{1}+r_{2})=2.5$. The results for our default setting are presented in Table 15. For completeness, we also include the results of training the regression model on merged 2-step experimental data with $c_{1}=0$ and varying $c_{t}$ in Table 16.
@@ 5.6 Dynamic Graph Structure
Moreover, we go back to use the classic linear potential outcome in Example 1 and examine the influence of dynamic graph structure. The main parameter is set as $(\beta_{0},\beta_{1},\beta_{2},\sigma_{e})=(1,1,1,0.1)$, as well as cases before. Even if the graph structure evolves, the GATE maintains to be $\beta_{1}+\beta_{2}=2$. We present the result in Table 17 and 18.
@@ A.1 Proof of Theorem 1
$\hat{\theta}_{1}=\beta_{1}+\beta_{2}\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{\top}B\mathbf{z}+\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{\top}\epsilon$ (57)
@@
$\mathbb{E}[\hat{\theta}_{1}]-(\beta_{1}+\beta_{2}\frac{1}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B))=\frac{\beta_{2}}{nc(1-c)}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}]-\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)$ (58)
@@
$\text{Bias}=-\frac{\beta_{2}}{n}(\sum_{i,j}B_{ij})\left(\frac{1}{n(n-1)}+1\right)$ (60)
@@
$\operatorname{Var}(\hat{\tau})=(\frac{\beta_{2}}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)+\frac{1}{\left(nc(1-c)\right)^{2}}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}(\mathbf{z}-c\mathbf{1})]\sigma^{2}_{e}$ (62)
@@
$(\frac{\beta_{2}}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)$ (64)
@@
$\left(\frac{\beta_{2}}{c(1-c)n}\right)^{2}\sum_{i,j,k,l}B_{ij}B_{kl}\mathbb{E}[(z_{i}-c)(z_{k}-c)z_{j}z_{l}]$ (65)
@@ A.2 Proof of Corollary 1
$\text{Bias}=-\frac{\beta_{2}}{n}(\sum_{i,j}B_{ij})\left(\frac{1}{n(n-1)}+1\right)=\Theta(1)$ (72)
@@
$\frac{\sigma_{e}^{2}}{\beta_{2}^{2}nc(1-c)}=\Theta(\frac{1}{n})$ (73)
@@ A.3 Proof of Theorem 2
$\phantom{=}\ \ \text{Bias}/(\beta_{2}\sum_{i,j}B_{ij})+1$ (89)
@@
$\operatorname{Bias}(\hat{\tau})/\beta_{2}=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{(n-1)\left(c_{1}-c_{2}\right)^{2}+2\left(c_{1}^{2}+c_{2}^{2}\right)-2\left(c_{1}+c_{2}\right)}{(n-1)\left(c_{1}+c_{2}\right)\left(2-\left(c_{1}+c_{2}\right)\right)}\right)$ (90)
@@ A.4 Proof of Theorem 3
$\text{Bias}/(\beta_{2}\sum_{i,j}B_{ij})+1$ $=\frac{Tn}{\sum_{t=1}^{T}d_{t}(Tn-\sum_{t=1}^{T}d_{t})}\left(\frac{\sum_{t=1}^{T}d_{t}(d_{t}-1)}{n(n-1)}-\frac{(\sum_{t=1}^{T}d_{t})^{2}}{Tn^{2}}\right)$ (91)
@@
$\operatorname{Bias}(\hat{\tau})/\beta_{2}$ $=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{T\sum_{t=1}^{T}c_{t}^{2}-(\sum_{t=1}^{T}c_{t})^{2}}{(\sum_{t=1}^{T}c_{t})(T-\sum_{t=1}^{T}c_{t})}\right)+O(\frac{1}{n})$ (92)
@@ A.7 Proof of Theorem 4
$\operatorname{Var}(\hat{\tau})/\beta_{2}^{2}=\Theta\left(\frac{1}{n}\right)$ (102)

### v3

@@ 5.2 Basic Simulation under Linear Interference
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+r\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\operatorname{deg}_{i}}+\epsilon_{i}$ (38)
where $\epsilon_{i}\sim N(0,\sigma_{e}^{2})$ are i.i.d. noises. We set the parameters as $(\beta_{0},\beta_{1},r,\sigma_{e})=(1,1,1,0.1)$, and the true GATE is $\beta_{1}+r=2$.
@@
Furthermore, if cluster-level randomization is employed and the ramp-up process has reached the stage of $c_{t}=0.05$, the bias is only $-0.032$, which is two orders of magnitude smaller than the signal scale, i.e., $\beta_{1}$ and $r$. Although the standard deviation remains $0.121$, which is still notable, the GATE estimate at this stage can be considered credible. Finally, for $c_{t}=0.5$, the performance reaches a state-of-the-art level compared to settings where only base levels and outcomes under $c_{T}$ are available, as far as we know.
@@ 5.3 Multiplicative Potential Outcome Model
$Y_{i}(\mathbf{z})=(\beta_{0}+\epsilon_{i})\cdot\frac{\deg_{i}}{\overline{\deg}}\cdot(1+\beta_{1}z_{i}+r\frac{\sum_{j\in\mathcal{N}_{i}}z_{j}}{\deg_{i}})$ (41)
Here, $\overline{\deg}$ is the average degree among units, and $\epsilon\sim N(0,\sigma_{e}^{2})$ are i.i.d. noises. Similar to previous section, we set $(\beta_{0},\beta_{1},r,\sigma_{e})=(1,1,1,0.1)$, and the true GATE is $\beta_{0}(\beta_{1}+r)=2$. We first present the result in Table 9.
@@ 5.4 Quadratic and Square Root Interference Term
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+r\left(\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\operatorname{deg}_{i}}\right)^{2}+\epsilon_{i}$ (42)
The quadratic interference term indicates that the marginal benefit of allocating more treatments increases constantly, making extrapolation more challenging. We set $(\beta_{0},\beta_{1},r,\sigma_{e})=(1,1,1,0.1)$, so the true GATE remains $\beta_{1}+r=2$. The results for our default setting are presented in Table 11, and the results with 2-step merged data, where $c_{1}=0$ and $c_{T}$ is varied, are shown in Table 12.
@@
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+r\sqrt{\left(\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\operatorname{deg}_{i}}\right)}+\epsilon_{i}$ (43)
with the same parameters $(\beta_{0},\beta_{1},r,\sigma_{e})=(1,1,1,0.1)$. We present corresponding results in Table 13 and 14.
@@ 5.5 Multihop interference
Moreover, we implement the multihop interference as shown in Example 2. Specifically, we maintain the setting that $(\beta_{0},\beta_{1},\sigma_{e})=(1,1,0.1)$, and set $M=2$ and $(r_{1},r_{2})=(1,0.5)$. Therefore, the true GATE in this case is given by:
$\text{GATE}=(\beta_{0}+\beta_{1}+\frac{1}{n}\mathbf{1}^{\top}B\mathbf{1})-\beta_{0}\approx 2.489$ (44)
The numerical outcome depends on the concrete network topology since we remove the circular effect that is introduced by the power of the (normalized) adjacency matrix. If we did not remove the diagonal terms of the interference matrix $B$, the GATE would be $\beta_{1}+(r_{1}+r_{2})=2.5$. The results for our default setting are presented in Table 15. For completeness, we also include the results of training the regression model on merged 2-step experimental data with $c_{1}=0$ and varying $c_{t}$ in Table 16.
@@ 5.6 Dynamic Graph Structure
Moreover, we go back to use the classic linear potential outcome in Example 1 and examine the influence of dynamic graph structure. The main parameter is set as $(\beta_{0},\beta_{1},r,\sigma_{e})=(1,1,1,0.1)$, as well as cases before. Even if the graph structure evolves, the GATE maintains to be $\beta_{1}+r=2$. We present the result in Table 17 and 18.
@@ A.1 Proof of Theorem 1
$\hat{\theta}_{1}=\beta_{1}+\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{\top}B\mathbf{z}+\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{\top}\epsilon$ (54)
@@
$\mathbb{E}[\hat{\theta}_{1}]-(\beta_{1}+\frac{1}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B))=\frac{1}{nc(1-c)}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}]-\frac{1}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)$ (55)
@@
$\text{Bias}=-\frac{1}{n}(\sum_{i,j}B_{ij})\left(\frac{1}{n(n-1)}+1\right)$ (57)
@@
$\operatorname{Var}(\hat{\tau})=(\frac{1}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)+\frac{1}{\left(nc(1-c)\right)^{2}}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}(\mathbf{z}-c\mathbf{1})]\sigma^{2}_{e}$ (59)
@@
$(\frac{1}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)$ (61)
@@
$\left(\frac{1}{c(1-c)n}\right)^{2}\sum_{i,j,k,l}B_{ij}B_{kl}\mathbb{E}[(z_{i}-c)(z_{k}-c)z_{j}z_{l}]$ (62)
@@ A.2 Proof of Corollary 1
$\text{Bias}=-\frac{1}{n}(\sum_{i,j}B_{ij})\left(\frac{1}{n(n-1)}+1\right)=\Theta(1)$ (69)
@@
$\frac{\sigma_{e}^{2}}{nc(1-c)}=\Theta(\frac{1}{n})$ (70)
@@ A.3 Proof of Theorem 2
$\phantom{=}\ \ \text{Bias}/(\sum_{i,j}B_{ij})+1$ (86)
@@
$\operatorname{Bias}(\hat{\tau})=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{(n-1)\left(c_{1}-c_{2}\right)^{2}+2\left(c_{1}^{2}+c_{2}^{2}\right)-2\left(c_{1}+c_{2}\right)}{(n-1)\left(c_{1}+c_{2}\right)\left(2-\left(c_{1}+c_{2}\right)\right)}\right)$ (87)
@@ A.4 Proof of Theorem 3
$\text{Bias}/(\sum_{i,j}B_{ij})+1$ $=\frac{Tn}{\sum_{t=1}^{T}d_{t}(Tn-\sum_{t=1}^{T}d_{t})}\left(\frac{\sum_{t=1}^{T}d_{t}(d_{t}-1)}{n(n-1)}-\frac{(\sum_{t=1}^{T}d_{t})^{2}}{Tn^{2}}\right)$ (88)
@@
$\operatorname{Bias}(\hat{\tau})$ $=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{T\sum_{t=1}^{T}c_{t}^{2}-(\sum_{t=1}^{T}c_{t})^{2}}{(\sum_{t=1}^{T}c_{t})(T-\sum_{t=1}^{T}c_{t})}\right)+O(\frac{1}{n})$ (89)
@@ A.7 Proof of Theorem 4
$\operatorname{Var}(\hat{\tau})=\Theta\left(\frac{1}{n}\right)$ (99)